# Data-Driven Probabilistic Air-Sea Flux Parameterization

### v2

@@ Abstract
Marked seasonal responses are obtained in deterministic and stochastic tests performed on forced single-column upper-ocean model GOTM.
@@ 3.1 Statistical scores of the deterministic predictions
Figure 2: (a) ANN-based deterministic predictions (orange dots) and bulk algorithm predictions (black dots) plotted against measured fluxes. We are only visualizing 10% of total samples. (b) $R^{2}$ and bias of ANN (orange) and bulk algorithm (black), evaluated on different geographical subsets. (c) ANN prediction of the mean of $Q_{S}$ on a uniform input grid. The dashed gray lines mark zero temperature difference $T_{a}$ = $T_{o}$. (d) ANN prediction of the std of $Q_{S}$ for the same grid. (e) Scatter plot of all data points and their marginal distribution. Gaussian kernel density estimation provides the gray contour lines that indicate the available samples per unit $\Delta U$ and $\Delta(T_{a}-T_{o})$. The same contour lines are overlaid on (c) and (d).
@@
The statistical scores of both ANN and COARE vary considerably between different turbulent fluxes. The sensible heat flux $Q_{S}$ is the hardest to predict for both based on its lowest $R^{2}$ values, while the latent heat flux $Q_{L}$ gives the largest magnitude of error in terms of RMSE. Another observation is that scores of both ANN and COARE also vary considerably between geographical locations. Figure 2(b) shows $R^{2}$ and bias of three fluxes evaluated on different subsets of cruises grouped roughly by geographical locations (same as in Figure 1). This is because each geographical subset has substantially different distributions of the input variable, as shown in Figure 1(c). In particular, $R^{2}$ for sensible heat flux $Q_{S}$ is very low in the Tropics, only around 0.15. This is partly because the variance of the sensible heat flux is small in the Tropics. Overall, the fluctuation of predictive scores over different regions is shared between ANN and COARE.
@@
The most salient improvement of ANN compared to baseline bulk algorithm is for latent heat flux, which is generally the least constrained of all turbulent fluxes. In certain regions, ANN improves $R^{2}$ up to about 0.2 increase in $R^{2}$. Overall, we notice that the estimate of latent heat given by bulk algorithm seems biased (towards predicting more evaporation from the ocean), and ANN fix the bias to some extent, as shown in Figure 2(b). This bias correction in latent heat flux was also found in a recent paper using data-driven method [zhou_physical-informed_2024] and may have important implications for global simulations. Supplementary table S1 lists RMSE and $R^{2}$ evaluated on the full dataset and on different geographical subsets. The reason underlying the statistical improvement is likely that the flexible structure of ANN provides more expressibility, and therefore fits the data better where certain assumptions of traditional bulk algorithms break down, e.g., the logarithmic profile for humidity and the expression for scalar roughness lengths. This idea is further discussed in the supplementary information and ANN predictions in terms of transfer coefficients $C_{D}$, $C_{H}$, and $C_{E}$ are also shown (Figure S1).
@@ 3.2 Structure of the mean and variance predicted by ANNs
In addition to the statistical scores, we further evaluate the ANN predictions on a reduced-dimension uniform grid to probe and interpret the ANN flux model. This is especially important for the std $\sigma$ as a way to understand how the predicted std depends on input variables. Figure 2(c) and (d) illustrate this using sensible heat flux as an example. Among the five inputs, we fix sea surface temperature $T_{o}$ = 10 $^{\circ}C$, relative humidity $RH=80\%$, and sea level pressure $p_{a}$ = 1010 hPa, while varying wind speed $U_{a}$ and temperature difference $T_{a}-T_{o}$. The plot represents a 2D slice of the 5D input space with typical values for the other variables.
@@
4 Tests: surface flux forced upper ocean mixing
As a preliminary test for the proposed probabilistic air-sea flux model, we choose to implement it in the single column model GOTM [umlauf_second-order_2005], which provides a controlled environment with no complex interplay with other nonlinear processes in GCMs. The governing equations are a set of 1D diffusion-type equations
@@ 4 Tests: surface flux forced upper ocean mixing
where the prognostic variables are the horizontal velocity components $u$ and $v$ (subject to Coriolis force), temperature $T$, and salinity $S$. The turbulent fluxes $\overline{w^{\prime}u^{\prime}}$, $\overline{w^{\prime}v^{\prime}}$, $\overline{w^{\prime}T^{\prime}}$, $\overline{w^{\prime}S^{\prime}}$ are not resolved but are instead parameterized by vertical mixing schemes. In particular, we test our air-sea flux algorithms in combination with two ocean surface boundary layer (OSBL) mixing parameterizations: the K-profile-parameterization (KPP) scheme [large_oceanic_1994] and a more sophisticated second-order closure $k-\epsilon$ scheme [umlauf_second-order_2005], both used in GCMs but typically under different resolutions.
Air-sea momentum and heat fluxes serve as boundary conditions to Equation 10. Additionally, surface fluxes can impact the mixing parameterization. In the KPP scheme, a change in heat flux (therefore buoyancy flux), affects the OSBL depth through the Richardson number criterion, and the sign of buoyancy flux determines the shape function [large_oceanic_1994]. In the $k-\epsilon$ scheme, the production terms in the turbulence kinetic energy (TKE) equation are affected by surface fluxes.
4.1 Heat flux time series at OWS Papa
@@
We use Ocean Weather Station (OWS) Papa (50°N, 145°W) as our test site. This long-term North Pacific station provides hourly meteorological and oceanic measurements, including temperature and salinity profiles. Its weak horizontal advection makes it a common benchmark for testing vertical mixing parameterization in single-column models [umlauf_second-order_2005, li_integrating_2021]. We discuss the flux predictions given the observed input variables at OWS Papa before showing results from the forced experiments.
@@
Figure 3(a) shows an example of ANN-predicted (deterministic) THF time series $Q=\mu_{Q_{S}}+\mu_{Q_{L}}$, compared to bulk algorithms, zoomed in to July and October 2015. We plot two bulk algorithms - COARE 3.6 and NCAR [large_global_2009] - to show that the ANN deviations exceed those between different bulk algorithms. This confirms that ANN is structurally different from any bulk algorithm. As with the THF itself, the difference between the ANN and bulk predictions is seasonal, as the input distributions vary throughout the year. This seasonality can be seen in Figure 3(c), which shows the monthly averaged flux discrepancy between ANN and COARE 3.6. The months of July and October see the most negative (more heat flux out of the ocean) and positive (less heat out) differences. Notice that within a period of high THF around Oct 10, ANN predicts the least extreme negative values while NCAR predicts the most extreme negative value, reflecting all three algorithms’ divergence at high wind speed, where few measurements exist to constrain them.
@@
Figure 3: (a) Example THF time series for July and October 2015, computed using observed input variables at OWS Papa. Black solid line: COARE 3.6; black dashed: NCAR; orange line: ANN. Orange shade shows one standard deviation for THF predicted by ANN, which is used to generate noise-perturbed heat fluxes (examples shown by thin orange lines). (b) Difference of temperature profiles in the deterministic single-column run $\Delta T(z,t)=T_{\text{ANN}}(z,t)-T_{\text{bulk}}(z,t)$. Black and orange dashed lines show the diagnosed MLD for bulk and ANN, respectively. (c) Monthly-averaged THF discrepancies between ANN and COARE. (d) SST and (e) MLD responses in the deterministic runs between (ANN minus COARE). Trends with inconsistent signs over the years are masked with hatching. (f) Monthly-averaged SST and (g) monthly-averaged MLD spread in 20 ensemble runs.
@@
Figure 3(a) also shows the ANN-predicted $\pm 1\>\sigma_{Q}$ for THF in orange shading. We assume no covariance between the sensible and latent heat flux residuals; therefore, the uncertainty of their sum simply follows $\sigma_{Q}^{2}=\sigma_{Q_{S}}^{2}+\sigma_{Q_{L}}^{2}$. Importantly, the predicted $\sigma_{Q}$ is state dependent, so the uncertainty envelope changes over time. Given the ANN-predicted time series of mean and std of THF, we can create an ensemble of stochastically perturbed fluxes. Two examples of such perturbed fluxes are shown with thin orange lines. The stochastic perturbation is expected to stem from certain unobserved processes with characteristic correlation time scales. Therefore, instead of white noise, we use temporally correlated noise $\epsilon$ generated through a discrete autoregressive model of order one - AR(1). Similar stochastic perturbation has been applied to momentum flux in the context of studying the diurnal cycle of ocean surface layer [giglio_role_2017] but not to heat flux on a seasonal to annual time scale, to our knowledge. The correlation time of 60 hours is estimated from the OWS Papa data (see supplementary information for details). It is also interesting to note that the bulk algorithm predicted THF is generally within $\pm\sigma_{Q}$, and thus can be interpreted as one possible ensemble member.
There are two parts of the probabilistic flux algorithms to test. First, we evaluate the effects of changing the flux algorithm from COARE 3.6 to ANN-predicted mean in deterministic runs. Second, we examine the uncertainty introduced by surface fluxes to the modeled upper-ocean state by performing ensemble runs using stochastically perturbed fluxes. Details of the numerical experiments are provided in the supplementary information. There is still a known net heat flux imbalance due to the omission of horizontal advection, which we mitigate by restarting the simulation with the observed profiles every month. For this and other reasons discussed in the supplementary information, our analysis emphasizes intercomparison between simulations rather than direct validation against observations, in particular two characteristic quantities: sea surface temperature (SST) and mixed layer depth (MLD).
4.2 Response in deterministic runs
@@
Figure 3(b) shows the difference in temperature profiles between ANN and COARE flux forced runs, as well as the MLD, for 2015. The composite statistics for four years of monthly-averaged change in SST and MLD are summarized in Figure 3(d) and (e). There is a perceivable response in temperature (fairly uniform across the mixed layer) as a result of a change in surface heat flux. ANN generally predicts more surface heat loss, which deepens the mixed layer and cools the sea surface. When the mixed layer is the shallowest in the summer months, there is the largest change in SST, consistent with a heat budget idea. We note that these effects are consistent across $k-\epsilon$ and KPP (see Figure S6), suggesting an insensitivity to OSBL parameterization.
In summary, switching from COARE 3.6 to ANN can result in differences of up to 2 meters in monthly-averaged MLD and 0.25 degrees in monthly-averaged SST. For context, switching the vertical mixing parameterization from $k-\epsilon$ to KPP can generate a much larger response in monthly averaged MLD, up to 18 meters (see Figure S4). Since perturbing the surface heat flux alone does not appear to generate a big response in MLD, this suggests that it is possible to calibrate surface heat flux and mixing separately, despite their interwoven effects on the upper ocean state. The response to surface heat flux is mainly in the spring and summer (at least at OWS Papa), while the response to mixing extends into the winter months. Future work will test perturbation to momentum flux, which might be dynamically more important.
4.3 Spread in stochastic ensemble runs
The ensemble of forced simulations provides an estimate of the variability in SST and MLD induced by the variability of THF. From the 20 ensemble member runs, we do not observe systematic drift induced by the stochastic forcing; the ensemble mean of the stochastic runs closely tracks the deterministic run (see Figure S5). This is expected since we are performing forced experiments and there is no feedback of state variables (such as SST) to surface forcing.
However, there is a certain level of ensemble spread, which is summarized in Figure 3(f) and (g). Here we focus on the spread in monthly averaged quantities, while the instantaneous spread can be larger (see Figure S5). The largest spread in MLD occurs during the months of April and May, since stochasticity can significantly affect the onset of restratification. This may have important implications for biogeochemical processes, suggesting that comparisons between modeled and observed upper ocean profiles, especially during restratification, may be sensitive to surface THF uncertainty. The largest spread in SST is seen around August, again because the mixed layer is shallow and SST is sensitive to surface heat fluxes. As in deterministic runs, the effects seen in stochastic runs are not particularly sensitive to different OSBL parameterizations.
@@ 5 Conclusion and discussion
We propose an ANN-based probabilistic model for turbulent air-sea fluxes. State-of-the-art bulk algorithms represent the statistical mean of eddy-covariance flux observations for a given set of state variables. Our work reinforces this idea by providing an alternative mean estimate using a purely data-driven approach. The ANN-predicted mean is similar to existing bulk algorithms, with marginally higher statistical correlation to data and reduced bias for latent heat flux. Additionally, the ANN-predicted standard deviation around the mean offers a measure of uncertainty (variability). Tests using OWS Papa data and a single-column model show that the difference in THF between the two algorithms’ estimates exhibits signs of seasonality, as does the response of upper ocean states, despite relatively small magnitudes. Stochastic ensemble runs indicate that the spread of MLD is largest during spring restratification, while the spread of SST is largest during summer when the mixed layer is shallow.
@@
We comment that the current framework can also be used to estimate the standard deviation around a given mean estimation, e.g., any existing bulk algorithm. In other words, we consider the residual between the observation and the bulk algorithm (or ANN) prediction. It is important to note that the uncertainty we estimate does not distinguish between measurement uncertainty (both instrument and sampling, with sampling uncertainty being the dominant factor) and missing physics (the inability to account for all predictive input variables). Such a partition is difficult and likely depends on the averaging window of the EC flux, which will be examined in future studies. Nonetheless, the ANN-estimated uncertainty highlights regions in the input space where the data is more “scattered”. This motivates field campaigns for better sampling of these regimes and examination of measurement uncertainty, as well as the use of high-fidelity numerical simulation to constrain the flux estimate [clayson_new_2023].
@@
There are several possible extensions of this work in the future. The ANN model uses a limited set of input variables, which is why the improvement over the existing algorithm may be considered marginal. Further improvement is possible, as several additional variables are measured and available in the dataset (e.g., turbulent kinetic energy). However, we must consider that many of these variables are not yet prognostic variables in GCMs. The height dependence is difficult to learn from the current data set alone, because there is little variance in the measurement height, but incorporating data from other measurement platforms (e.g., buoy) might help. We also need more high-quality vertical profile measurements of surface wind, temperature, and humidity. The probabilistic model presented assumes a univariate conditional Gaussian distribution for each flux component. This may not fully capture the underlying complexity, especially if the residual is induced by unrepresented physical processes. The model can be extended to include the covariance between flux components, and the Gaussian assumption can be relaxed, given that we have more data to train a more complicated statistical model. Finally, the analysis at OWS Papa provides an initial estimate of the response magnitude but is unlikely to be representative of the global ocean with varying meteo-ocean conditions and coupling dynamics across regions. We plan to extend our analysis to global simulations to assess the broader impact of modified air-sea flux algorithm.
@@
While existing bulk algorithms represent the mean values of turbulent fluxes given limited input variables, our approach takes a step towards examining the variability around the mean values. Stochastic air-sea flux parameterization offers a promising alternative to the deterministic approach. It is crucial to prescribe the appropriate magnitude and correlation scale of noise with varying model resolutions, which needs to be better understood in future studies. In the single-column test, the spread induced by stochastic residuals exceeds that of different deterministic flux algorithms. Therefore, extending the testing to global GCMs could potentially affect the variability and address long-standing biases through interaction with horizontal transport and other nonlinear dynamics.
@@
The full NOAA PSL cruises are documented and available here https://downloads.psl.noaa.gov/psd3/cruises/. For this study we used a compact compilation data set https://github.com/jiarong-wu/mlflux/blob/main/fluxes_all_cruises_compilation.nc. The code for ANN training and evaluation can be found in this repository https://github.com/jiarong-wu/mlflux. The GOTM code is available here https://gotm.net/portfolio/. In particular, this work used an implementation of GOTM in the PDE solver Basilsik [popinet_vertically-lagrangian_2020] http://basilisk.fr/src/test/ows_papa.c. For comparison to bulk algorithms, we used the aerobulk-python package https://github.com/xgcm/aerobulk-python.

### v3

@@ Abstract
Marked seasonal sensitivities are found in deterministic and stochastic tests performed on forced single-column upper-ocean model GOTM.
Plain Language Summary
Understanding the exchange of heat and momentum between the ocean and the atmosphere is key to improving weather and climate predictions. However, quantifying these exchanges (fluxes) is difficult, and models rely on simplified equations to compute the flux values based on state variables (wind speed, air temperature, sea surface temperature, etc.). This study introduces a new statistical method based on machine learning (artificial neural networks) trained using direct field observations of fluxes to better estimate these ocean-atmosphere fluxes. The model provides not only a better estimate of the average flux, but also an understanding of its uncertainty. Preliminary tests show that this new flux estimation method has a considerable effect on the simulated state of the upper ocean, especially during certain seasonal shifts. This approach helps improve the accuracy of air-sea flux estimates and can eventually lead to better coupled weather and climate models.
@@ 3.1 Statistical scores of the deterministic predictions
Figure 2: (a) ANN-based deterministic predictions (orange dots) and bulk algorithm predictions (black dots) plotted against measured fluxes. We are only visualizing 10% of total samples. (b) $R^{2}$ and bias of ANN (orange) and bulk algorithm (black), evaluated on different geographical subsets. (c) ANN prediction of the mean of $Q_{S}$ on a uniform input grid. Among the five inputs, we fix sea surface temperature $T_{o}$ = 10 $^{\circ}C$, relative humidity $RH=80\%$, and sea level pressure $p_{a}$ = 1010 hPa. The dashed gray lines mark zero temperature difference $T_{a}$ = $T_{o}$. (d) ANN prediction of the std of $Q_{S}$ for the same grid. (e) Scatter plot of all data points and their marginal distribution. Gaussian kernel density estimation provides the gray contour lines that indicate the available samples per unit $\Delta U$ and $\Delta(T_{a}-T_{o})$. The same contour lines are overlaid on (c,d).
@@
The statistical scores of both ANN and COARE vary considerably between different turbulent fluxes. The sensible heat flux $Q_{S}$ is the hardest to predict for both based on its lowest $R^{2}$ values, while the latent heat flux $Q_{L}$ gives the largest magnitude of error in terms of RMSE. Another observation is that scores of both ANN and COARE also vary considerably between geographical locations. Figure 2(b) shows $R^{2}$ and bias of three fluxes evaluated on different subsets of cruises grouped roughly by geographical locations (same as in Figure 1). This is because each geographical subset has substantially different distributions of the input variable, as shown in Figure 1(c). In particular, $R^{2}$ for sensible heat flux $Q_{S}$ is very low in the Tropics, only around 0.15. This is partly because the variance of the sensible heat flux is small there. Overall, the fluctuation of predictive scores over different regions is shared between ANN and COARE.
@@
The most salient improvement of ANN compared to baseline bulk algorithm is for latent heat flux, which is generally the least constrained of all turbulent fluxes. In certain regions, ANN sees up to around 0.2 increase in $R^{2}$. Overall, we notice that the estimate of latent heat given by bulk algorithm seems biased (towards predicting more evaporation from the ocean), and ANN fixes the bias to some extent, as shown in Figure 2(b). This bias correction in latent heat flux was also found in a recent paper using data-driven method [zhou_physical-informed_2024] and may have important implications for global simulations. We comment that a latent heat flux model accounting for sea spray generation may offer more flexibility and help mitigate this bias [andreas_improved_2015, barr_sea-state-dependent_2023]. Supplementary table S1 lists RMSE and $R^{2}$ evaluated on the full dataset and on different geographical subsets. The reason underlying the statistical improvement is likely that the flexible structure of ANN provides more expressibility, and therefore fits the data better where certain assumptions of traditional bulk algorithms break down, e.g., the logarithmic profile for humidity and the expression for scalar roughness lengths. This idea is further discussed in the supplementary information and ANN predictions in terms of transfer coefficients $C_{D}$, $C_{H}$, and $C_{E}$ are also shown (Figure S1).
@@ 3.2 Structure of the mean and variance predicted by ANNs
In addition to the statistical scores, we further evaluate the ANN predictions on a reduced-dimensional uniform grid to probe and interpret the ANN flux model. This is especially important for understanding how the predicted std depends on input variables. Figure 2(c) and (d) illustrate this using sensible heat flux as an example.
@@
4 Idealized sensitivity tests of probabilistic air-sea flux parameterization
To assess the impact of the proposed stochastic air–sea flux parameterization, we implement it in the single column model GOTM [umlauf_second-order_2005], which provides a controlled environment without the complex interactions with other nonlinear processes present in GCMs and significantly reduces the computational cost of running ensembles. The governing equations are a set of 1D diffusion-type equations
@@ 4 Idealized sensitivity tests of probabilistic air-sea flux parameterization
where the prognostic variables are the horizontal velocity components $u$ and $v$ (subject to Coriolis force), temperature $T$, and salinity $S$. The turbulent fluxes $\overline{w^{\prime}u^{\prime}}$, $\overline{w^{\prime}v^{\prime}}$, $\overline{w^{\prime}T^{\prime}}$, $\overline{w^{\prime}S^{\prime}}$ are not resolved but are instead parameterized by vertical mixing schemes. For robustness, we test our air-sea flux algorithms in combination with two ocean surface boundary layer (OSBL) mixing parameterizations: the K-profile-parameterization (KPP) scheme [large_oceanic_1994] and a more sophisticated second-order closure $k-\epsilon$ scheme [umlauf_second-order_2005], both used in GCMs but typically under different resolutions. Air-sea momentum and heat fluxes serve as boundary conditions to Equation 10. Additionally, surface fluxes can impact the mixing parameterization.
4.1 THF time series at OWS Papa
@@
We use Ocean Weather Station (OWS) Papa (50°N, 145°W) as our test site. This long-term North Pacific station provides hourly meteorological and oceanic measurements, including temperature and salinity profiles. Its weak horizontal advection makes it a common benchmark for testing vertical mixing parameterization in single-column models [umlauf_second-order_2005, li_integrating_2021]. Here, our goal is simply to use this location to demonstrate the use of stochastic flux parameterization and to assess the potential impact on upper ocean state, with given vertical mixing parameterizations. We discuss the flux predictions given the observed input variables at OWS Papa before showing results from the forced experiments.
@@
Figure 3(a) shows an example of ANN-predicted (deterministic) THF time series $Q=\mu_{Q_{S}}+\mu_{Q_{L}}$, compared to bulk algorithms, zoomed in to July and October 2015. We plot two bulk algorithms - COARE 3.6 and NCAR [large_global_2009] - to show that the ANN deviations exceed those between different bulk algorithms. As with the THF itself, the difference between the ANN and bulk predictions is seasonal, since the input distributions vary throughout the year. Figure 3(c) shows the monthly averaged flux discrepancy between ANN and COARE 3.6. The months of July and October see the most negative (more heat flux out of the ocean) and positive differences. Notice that within a period of high THF around Oct 10, ANN predicts the least extreme negative values while NCAR predicts the most extreme negative value, reflecting all three algorithms’ divergence at high wind speed, where few measurements exist to constrain them.
@@
Figure 3: (a) Example THF time series for July and October 2015, computed using observed input variables at OWS Papa. Black solid line: COARE 3.6; black dashed: NCAR; orange line: ANN. Orange shade shows one standard deviation for THF predicted by ANN, which is used to generate noise-perturbed heat fluxes (examples shown by thin orange lines). (b) Difference of temperature profiles in the deterministic single-column run $\Delta T(z,t)=T_{\text{ANN}}(z,t)-T_{\text{bulk}}(z,t)$. Black and orange dashed lines show the diagnosed MLD for bulk and ANN, respectively. (c) Monthly-averaged THF discrepancies between ANN and COARE. (d) SST and (e) MLD responses in the deterministic runs between (ANN minus COARE). Trends with inconsistent signs over the years are masked with hatching. (f) Monthly-averaged SST and (g) monthly-averaged MLD spread in 20 ensemble runs. The spread is monthly averaged while the instantaneous value can be larger (see Figure S5).
@@
Figure 3(a) also shows the ANN-predicted $\pm 1\>\sigma_{Q}$ for THF. We assume no covariance between the sensible and latent heat flux residuals, so the uncertainty of their sum simply follows $\sigma_{Q}^{2}=\sigma_{Q_{S}}^{2}+\sigma_{Q_{L}}^{2}$. Importantly, the predicted $\sigma_{Q}$ is state dependent, so the uncertainty envelope changes over time. Given the ANN-predicted mean and std, we can create an ensemble of stochastically perturbed fluxes, and two examples are shown in Figure 3(a). The stochastic perturbation is expected to stem from certain unobserved processes with characteristic correlation time scales. Therefore, instead of white noise, we use temporally correlated noise $\epsilon$ generated through a discrete autoregressive model of order one - AR(1) [giglio_role_2017]. The correlation time of 60 hours is estimated from the OWS Papa data. Note that the bulk algorithm predicted THF is generally within $\pm\sigma_{Q}$, and thus can be interpreted as one possible ensemble member.
Despite relatively weak horizontal advection, there is still a net heat flux imbalance at OWS Papa [large_oceanic_1994] if horizontal advection is ignored. We mitigate this by restarting the simulation with observed temperature and salinity profiles every month and by emphasizing inter-comparison between simulations rather than direct validation against observations. Details of the numerical experiments are provided in the supplementary information.
4.2 Sensitivity of the upper ocean state to THF uncertainty
With the monthly restart and stochastically perturbed THF ensemble, our tests can be understood as experiments that quantify the short-term uncertainty in the upper ocean state due to uncertainty in the flux algorithm. In the following discussion, we focus on two characteristic quantities: sea surface temperature (SST) and mixed layer depth (MLD).
@@ 4.2 Sensitivity of the upper ocean state to THF uncertainty
Related to this annual cycle, our main finding is a significant seasonality in the sensitivity to surface heat flux. Figure 3(b) shows the difference in temperature profiles between ANN and COARE flux forced runs, as well as the MLD, for 2015. There is a perceivable difference in temperature (fairly uniform across the mixed layer) as a result of a change in surface heat flux. ANN generally predicts more surface heat loss, which deepens the mixed layer and cools the sea surface. When the mixed layer is the shallowest in the summer months, there is the largest change in SST, consistent with a heat budget idea. The seasonality is confirmed by the composite statistics for four years of monthly-averaged change in SST and MLD are summarized in Figure 3(d) and (e). We note that these effects are consistent across $k-\epsilon$ and KPP (see Figure S6), suggesting that the effects of air-sea fluxes is somewhat insensitive to the OSBL parameterization used.
Perhaps more important than comparison between any two specific deterministic runs, the ensemble of stochastically perturbed simulations provides an estimate of the variability in SST and MLD induced by the variability of THF. From the 20 ensemble member runs, we do not observe systematic drift induced by the stochastic forcing; the ensemble mean of the stochastic runs closely tracks the deterministic run (see Figure S5). This is expected since we are performing forced experiments and there is no feedback of state variables (such as SST) to surface forcing. However, we observe considerable spread between ensemble members, which is summarized in Figure 3(f) and (g), again with clear seasonal patterns. The largest spread in MLD occurs during the months of April and May, since stochasticity can significantly affect the onset of restratification. This may have important implications for biogeochemical processes, suggesting that comparisons between modeled and observed upper ocean profiles, especially during restratification, may be sensitive to surface THF uncertainty. The largest spread in SST is seen around August, again because the mixed layer is shallow and SST is sensitive to surface heat fluxes. As in deterministic runs, the effects seen in stochastic runs are not particularly sensitive to different OSBL parameterizations.
@@ 5 Conclusion and discussion
We propose an ANN-based probabilistic model for turbulent air-sea fluxes. State-of-the-art bulk algorithms represent the statistical mean of eddy-covariance flux observations for a given set of state variables. Our work reinforces this idea by providing an alternative mean estimate using a purely data-driven approach. The ANN-predicted mean is similar to existing bulk algorithms, with marginally higher statistical correlation to data and reduced bias for latent heat flux. Additionally, the ANN-predicted standard deviation around the mean offers a measure of uncertainty (variability). Tests using OWS Papa data and a single-column model show that the difference in THF between the different algorithms’ estimates exhibits signs of seasonality, as does the sensitivity of upper ocean states, despite relatively small magnitudes. Stochastic ensemble runs indicate that the spread of MLD is largest during spring restratification, while the spread of SST is largest during summer when the mixed layer is shallow.
@@
We comment that the current framework can also be used to estimate the standard deviation around a given mean estimation, e.g., any existing bulk algorithm, by considering the residual between the observation and the mean estimation. It is important to note that the uncertainty we estimate does not distinguish between measurement uncertainty [<]both instrument and sampling, with sampling uncertainty being the dominant factor, see for example¿rannik_random_2016 and missing physics (the inability to account for all predictive input variables). Such a partition is difficult and likely depends on the averaging window of the EC flux, which will be examined in future studies. Nonetheless, the ANN-estimated uncertainty highlights regions in the input space where the data is more “scattered”. This motivates field campaigns for better sampling of these regimes and examination of measurement uncertainty, as well as the use of high-fidelity numerical simulation to constrain the flux estimate [clayson_new_2023].
@@
There are several possible extensions of this work in the future. The ANN model uses a limited set of input variables, which is why the improvement over the existing algorithm may be considered marginal. Further improvement is possible, as several additional variables are measured and available in the dataset (e.g., turbulent kinetic energy), but many of them are not yet prognostic variables in GCMs. The height dependence is difficult to learn from the current data set alone, but incorporating data from other measurement platforms (e.g., buoy) might help. We also need more high-quality vertical profile measurements of surface wind, temperature, and humidity. The probabilistic model presented assumes a univariate conditional Gaussian distribution for each flux component, which may not fully capture the underlying complexity of unobserved physical processes. The model can be extended to relax the Gaussian assumption and include the covariance between flux components. It would also be valuable to explore potential connections between the data-driven formulation and ad hoc modifications commonly used in flux algorithms, such as the gustiness factor. Finally, the analysis at OWS Papa provides an initial estimate of the sensitivity magnitude but is unlikely to be representative of the global ocean with varying meteo-ocean conditions and coupling dynamics across regions. We plan to extend our analysis to global simulations to assess the broader impact of the proposed air-sea flux algorithm.
@@
In summary, the probabilistic framework proposed in this paper provides a formal way to represent uncertainty in air–sea flux estimates, which is particularly important given the extreme scarcity of direct observations, despite their continued role as the primary means for constraining air–sea flux parameterizations. While existing bulk algorithms represent the mean values of turbulent fluxes given limited input variables, our approach takes a step towards examining the variability around the mean values. Stochastic air-sea flux parameterization offers a promising alternative to the deterministic approach. It is crucial to prescribe the appropriate magnitude and correlation scale of noise with varying model resolutions, which needs to be better understood in future studies. In the single-column test, the spread induced by stochastic residuals exceeds that of different deterministic flux algorithms. Therefore, extending the testing to global GCMs could potentially affect the variability and address long-standing biases through interaction with horizontal transport and other nonlinear dynamics.
Conflict of Interest
The authors declare no conflicts of interest relevant to this study.
@@
The NOAA PSL data are documented and available here https://downloads.psl.noaa.gov/psd3/cruises/. The code for ANN training and evaluation, as well as the model weights are openly available [jiarong_wu_2025_17554656]. The GOTM code is available here https://gotm.net/portfolio/. In particular, this work used an implementation of GOTM in the PDE solver Basilsik [popinet_vertically-lagrangian_2020] http://basilisk.fr/src/test/ows_papa.c. For comparison to bulk algorithms, we used the aerobulk-python package https://github.com/xgcm/aerobulk-python.